### v1

# Event2Vec: Processing neuromorphic events directly by representations in vector space

Abstract
The neuromorphic event cameras have overwhelming advantages in temporal resolution, power efficiency, and dynamic range compared to traditional cameras. However, the event cameras output asynchronous, sparse, and irregular events, which are not compatible with mainstream computer vision and deep learning methods. Various methods have been proposed to solve this issue but at the cost of long preprocessing procedures, losing temporal resolutions, or being incompatible with massively parallel computation. Inspired by the great success of the word to vector, we summarize the similarities between words and events, then propose the first event to vector (event2vec) representation. We validate event2vec on classifying the ASL-DVS dataset, showing impressive parameter efficiency, accuracy, and speed than previous graph/image/voxel-based representations. Beyond task performance, the most attractive advantage of event2vec is that it aligns events to the domain of natural language processing, showing the promising prospect of integrating events into large language and multimodal models. Our codes, models, and training logs are available at https://github.com/fangwei123456/event2vec.
keywords:
neuromorphic computing, event-based vision, brain-inspired computing
1 Introduction
Neuromorphic computing is an emerging research area that aims to build the next generation of artificial intelligence by taking inspiration from the brain [1]. One of the most significant advancements of this paradigm is the event camera [2], such as the Dynamic Vision Sensor (DVS) [3] and the Asynchronous Time-based Image Sensor (ATIS) [4]. Compared to the traditional cameras which output synchronous frames, the event cameras work in an asynchronous fashion and output events triggered by the change of brightness, showing extremely high temporal resolution, low power consumption, and High Dynamic Range (HDR) imaging ability.
Events are formulated in Address-Event-Representation (AER). The $i$-th event in an event stream is $E_{i}=(x_{i},y_{i},t_{i},p_{i})$, where $x_{i},y_{i}$ are the coordinates, $t_{i}$ is the timestamp and $p_{i}$ is the polarity. The range of $x_{i},y_{i}$ is determined by the sensor’s spatial resolution, e.g., $0\leq x_{i}\leq 127,0\leq y_{i}\leq 127$ for the DVS 128 camera. The temporal resolution of $t_{i}$ is in $\mu s$ range. $p_{i}\in\{0,1\}$ represents whether the brightness at position $x_{i},y_{i}$ is decreasing or increasing at timestamp $t_{i}$. The AER format is a sparse and irregular representation. However, most modern Computer Vision (CV) methods are based on dense and regular multi-dimensional tensor representation. For example, an image is represented as a tensor with the shape of $C\times H\times W$, where $C$ is the number of channels, $H$ is the height and $W$ is the width. Typically, $C=1$ for grayscale image and $C=3$ for color image. The multi-dimensional tensor representation is compatible with deep learning methods [5], modern machine learning, and scientific computing frameworks [6, 7, 8].
To apply powerful deep learning methods in event-based vision, there are two successful solutions. The first category is a conversion method that converts events to dense and regular multi-dimensional tensor representations. Time surface [9], event-to-frame [10], and event-to-voxel [11] methods belong to the conversion category. After the conversion, dense and regular tensors are obtained and can be used by modern deep learning methods. However, these methods diminish and/or destroy temporal resolution of events since timestamps are quantized during conversion, and hence, conversion imposes large latency in real-time tasks [12, 2]. The second category is processing asynchronous and sparse events directly. Spiking neural networks (SNNs) [13, 14], sparse convolutional networks (sparse CNNs) [15], and graph neural networks (GNNs) [16, 17] are typical network structures that inherently possess sparse and asynchronous computations. However, modern deep learning methods rely heavily on Graphics Processing Units (GPUs), which are not well-suited for dynamic computations and unstructured sparse acceleration. Thus, these methods cannot fully utilize the massively parallel computational ability of GPUs and are much slower than traditional CNNs or dense models.
2 View of Encoding
Existing methods can also be regarded as solving the encoding problem of events, i.e., extracting and converting information from events to other formats. Conversion methods encode events to tensors with specific physical significances. For example, frames integrated from events are the sum of brightness changes during the temporal integration interval. SNNs encode events to spikes with spiking neurons. Sparse CNNs regard events as sparse images. GNNs take events as graphs, where, event-to-graph pre-processing is also required.
The most popular encoding method in machine learning is the word to vector (word2vec) [18], which solves the issue of representing words. It embeds each word to a fixed-length vector (token), and then the relationship between different words can be expressed by the math operations between vectors. This method is fully compatible with deep learning methods and has achieved significant success in Natural Language Processing (NLP) tasks [19, 20]. Another typical example is the Vision Transformer (ViT) [21], which yields significantly higher accuracy that CNNs owing to attention-based Transformer structure [22]. A vital component of ViT is the patchifying encoding method, which embeds image grids to tokens and enables Transformers to process images directly.
There are many similarities between words and events. Fig.1 shows the comparison between them. Our conclusions are as follows:
  1. (1)
    Every element is a combination of an index and a position. Each word has a unique index in the dictionary. Converting words to indices is implemented by the tokenizer in NLP tasks, and the indices in Fig.1 are generated by the LLama-3 [23] tokenizer. The position of a word is determined by the location in the sentence, e.g., 00 for ”I” in the sentence ”I am glad to see you”. The index of an event is the tuple $(x_{i},y_{i},p_{i})$. However, the position is not $i$ but $t_{i}$, which indicates the temporal location in the event stream.
  2. (2)
    The number of indices is finite. A dictionary contains a certain number of words. An event camera can only output events with limited indices, e.g., $2\times 128\times 128$ for the DVS 128 camera, where $2$ is the number of polarities and $128\times 128$ is the spatial resolution.
  3. (3)
    The meaning of an element is determined by its context. A word may have multiple meanings. For example, ”transformer” means the neural network structure in deep learning, but can also represent roles in animated series. The meaning of a word is inferred by its context. An individual event indicates the brightness at a certain spatial and temporal location, but no more information can be inferred. However, a stream of events can form the outline of an object, then an event can be regarded as part of an edge. Thus, the meaning of an event is also determined by its context, the event stream.
Figure 1: Comparison between words and events. The figure of the DVS 128 camera is cited from [3].
3 Represent Event in Vector Spase
3.1 Event to Vector
Given the similarities between words and events, it is natural to represent events in vector space, i.e., event to vector (event2vec). For an event $(x,y,t,p)$ obtained from an event camera with $H\times W$ spatial resolution, we embed it to the unique embedding 1D-index $i$ and embedding vector $\textbf{v}\in\mathbb{R}^{d}$:
$i$ $=p\cdot H\cdot W+y\cdot W+x,$ (1)
v $=\text{Embed}(i),$ (2)
where $\text{Embed}(\cdot)$ is the embedding layer with learnable weights $\textbf{W}\in\mathbb{R}^{(2\cdot H\cdot W)\times d}$, and $d$ is the embedding dimension.
A key difference between words and events is their spatial structure. The word is embedded in a 1D flattened sentence, while the event occurs in a 2D plane. However, the naive event2vec, Eq.(1)-(2) does not consider this feature. Thus, we modify the naive event2vec to the 2D event2vec formulation:
$i$ $=p\cdot H+y,$ (3)
$j$ $=p\cdot W+x,$ (4)
v $=[\text{Embed}_{y}(i),\text{Embed}_{x}(j)].$ (5)
Eq.(3) and Eq.(4) first encode $y$ and $x$ to unique indices, then Eq.(5) concatenates the embedding tensors.
3.2 Neighbor Semantics
It is worth noting that the real physical world is continuous, but the pixel coordinates captured by a camera are discrete. The generation of $(x,y)$ for an event introduces noise and quantization. Say, when capturing an identical object twice, we may get $(x,y)$ the first time and $(x+1,y)$, $(x,y-1)$ or other neighbor coordinates the second time. Thus, we expect that events with close coordinates should have similar semantics. Compared to naive event2vec Eq.(1)-(2) which can only output the same v when both $x$ and $y$ are the same for two events, the 2D event2vec Eq.(3)-(5) will output the same $\textbf{v}_{y}$ or $\textbf{v}_{x}$ when input events have either same $y$ or $x$. However, it still does not fully account for the similarity between neighbors, such as $(x,y)$ and $(x-1,y-1)$. Although neural networks have the potential to learn this spatial-semantic similarity without any inductive bias, learning may be difficult, i.e., may require many training epochs and large datasets.
To achieve neighbor similarity, a trivial idea is downsampling the spatial resolution. For example, if two events have the coordinates $(0,0)$ and $(1,1)$ at $128\times 128$ resolution, their coordinates will be both $(0,0)$ after downsampling the resolution to $64\times 64$. In fact, this method is similar to patchifying in ViT which uses grids rather than pixels as the element.
The second solution is involving the neighbors of an event manually. When processing $(x,y)$, we regard it as a group of coordinates and embed both the central point $(x,y)$ and its neighbors. Denote the number of neighbors as $n$, we expand the 2D event2vec as:
$\textbf{v}_{y}$ $=\sum_{k=-n}^{k=n}w_{k}\cdot\text{Embed}_{y}(p\cdot H+y+k),$ (6)
$\textbf{v}_{x}$ $=\sum_{k=-n}^{k=n}w_{k}\cdot\text{Embed}_{x}(p\cdot W+x+k),$ (7)
v $=[\textbf{v}_{y},\textbf{v}_{x}],$ (8)
where $w_{k}$ is a learnable parameter and initialized by the distance:
$w_{k}=\exp\Big{(}\frac{-k^{2}}{2}\Big{)}.$ (9)
With this neighbor embedding, the semantics of an event is determined not only by itself but also by its neighbors. This embedding is more consistent with the 2D spatial structure of the events.
3.3 Temporal Positional Encoding
Compared with words, events have explicit positions, i.e. timestamps. Transformers use sinusoidal or rotational positional encoding methods to encode position information[22, 24]. To incorporate temporal information from events, we can also use temporal positional encoding. Positional encoding in NLP tasks can be regarded as encoding the positions of words $\{0,1,...,L-1\}$, where $L$ is the sequence length. For events, the only thing to do is replace the inputs by $\{t_{0},t_{1},...,t_{L-1}\}$. Additionally, timestamps can also be added to embedding vectors directly because they already have explicit physical semantics, i.e. the physical time of events occurring.
4 Experiments
We validate the proposed event2vec methods on the neuromorphic ASL-DVS [16] classification task. The ASL-DVS dataset contains 24 letters (classes) from the American Sign Language. Each class contains 4200 samples and each sample lasts about 100 $ms$. We use the first 80% of the dataset as the train set and the last 20% as the test set. The train-test ratio is same as [16]. To avoid any challenges with code reproducibility and allow fair comparison for researchers using our methodology, we do not split randomly.
4.1 Training Hyparameters
We use a vanilla transformer encoder structure with only two layers, 256 features, 16-head self-attention, and 512 features in the feedforward layers. Note that the space complexity of self-attention is $\mathcal{O}(L^{2})$, where $L$ is the sequence length. Thus, we cannot process event streams with too many events. We just randomly sample $L=255$ events from each sample. The sampling is implemented by the choice function of Numpy [8], and the randomness is controlled by the random generator in Numpy. Although the random sampling will cause undetermined test accuracy, we find that the results are stable after about averaging over four repeated tests, which is also the test accuracy reported in this article. We set all random seeds as 0. We use the AdamW [25] optimizer with a learning rate of 0.001, batch size = 128, and cosine annealing learning rate scheduler [26]. We train the model for 64 epochs. We downsample the spatial size from $180\times 240$ to $32\times 48$.
We directly add the timestamps to the embedding vectors and do not use any additional temporal positional encoding techniques. We find that this simple encoding method works well enough. We normalize the timestamps because their values are too large and may cause numerical instability in neural networks. The maximum value of the timestamp in the train set is $t_{max}=521217$. For both train and test sets, we normalize the timestamp by $t=\frac{t}{t_{max}}$. We regard that the absolute values of timestamps are meaningless, and we set all timestamps in an event stream minus the first timestamp.
The experiments are carried out to show the effectiveness of our proposed event2vec, rather than seeking state-of-the-art accuracy. Thus, we do not use any data augmentation or regularization such as dropout and weight decay.
4.2 Accuracy, Efficiency, and Speed
With $n=2$ neighbors and 2D event2vec embedding, our model achieves 99.68% test accuracy. Due to random sampling, the model only utilizes about 0.00143% of all events in the training set in one iteration. After 64 epochs (40319 iterations), only 58% of the train set is used to train the model. The high test accuracy shows the extreme learning efficiency of the model equipped with our event2vec representation. Tab.1 compares the event2vec representation with some previous representations on the ASD-DVS dataset. It is worth noting that our method achieves high accuracy with an extremely small model size and does not require any event data preprocessing. Our model achieves 2185 samples/s training speed and 4320 samples/s inference speed on a Red Hat Enterprise Linux 8 server with Intel(R) Xeon(R) Gold 6240 CPU @ 2.60GHz, a single Nvidia Tesla V100-SXM2-32GB GPU, and 64GB RAM. For comparison, [27] takes 16.7 $ms$ to process one sample, and the execution time does not include the event to frame and voxel-graph preprocessing. Our method only requires 1000/4320=0.23 $ms$ for one sample, which is about 72$\times$ faster than [27]. Note that our speed includes the data reading and sampling.
Fig.2 compares the test accuracy of different event2vec methods. Note that we transform the training iteration steps to the ratio of used training data for ease of reading. The results show that the spatial structure of the 2D event2vec is vital, without which the model can hardly learn the relationship between close coordinates, leading to much lower accuracy. With the weighted neighbor embedding, the accuracy of 2D event2vec is further promoted, but the accuracy saturates with more than 2 neighbors. Thus, using $n=2$ neighbors is sufficient.
Fig.3 shows the test accuracy under different downsampling sizes and event2vec methods. When using the naive event2vec, lower spatial resolution is helpful for clustering neighbor events and leads to higher accuracy. When using the 2D event2vec with weighted neighbors, the neighbor semantics are inherent, and the quantization of coordinates under low spatial resolution may cause accuracy loss, e.g., accuracy under $9\times 12$ resolution is slightly lower than that under $36\times 48$ when using 2D event2vec, but higher when using the naive event2vec.
Method Preprocessing Accuracy (%) Model Size (MB) Epoch
GNN+CNN [16] To graphs 90.1 19.46 150
GNN + Transformer [27] To images and voxel-graph 99.6 220.3 150
Event2vec + Transformer (Ours) None 99.68 4.13 64
Table 1: Comparison with previous representations on the ASL-DVS dataset.
(a) Effectiveness of 2D embedding
(b) Effectiveness of neighbors embedding
Figure 2: Effectiveness event2vec methods. (a) The spatial structure of the 2D event2vec is vital. (b) The induction of neighbors is helpful, but the effectiveness is marginal. In general, $n=2$ neighbors are good enough.
(a) Naive event2vec
(b) 2D event2vec with 2 neighbors.
Figure 3: Comparison of downsampling sizes. (a) Downsampling is helpful for the naive event2vec. (b) But its effectiveness is marginal, even harmful for 2D event2vec with neighbors.
5 Discussion
The neuromorphic event cameras bring new opportunities and challenges for computer vision. Researchers have made massive efforts to coordinate the gap between event-based vision and deep learning methods. In this article, we propose the first-of-its-kind event2vec representation and enable neural networks to process asynchronous events directly without losing any temporal resolution.
With the proposed event2vec representation, a vanilla transformer encoder with only 2 layers achieves impressive performance on the ASL-DVS classification task. The extremely small model size and high inference speed are more attractive than the test accuracy results. The parameter efficiency shows a promising prospect of deploying the model with event2vec on resource-limited edge computing devices. Although neuromorphic event cameras have $\mu s$-level temporal sensitivity, the computational speed of subsequent models cannot keep up with the preceding sensors in most cases. With the event2vec representation, the model gets rid of the latency of pre-processing the event data and has the potential to achieve $ms$-level response speed in an end-to-end scenario.
Beyond parameter efficiency, task accuracy, and running speed, the most promising feature of event2vec is that it aligns events to the domain of natural language processing, which has achieved significant progress and attracted interest even outside academia and industry. By combining event2vec representation with powerful large language models, various applications can be explored. For example, we can use a decoder-only transformer structure and train on events with self-supervised learning, then the model is expected to generate future events from previous events like an ”event-GPT”. We can also build multi-modal large models incorporating events, images, and audio for low-latency and robust autonomous driving under complex environments. In conclusion, our research has the potential to extend the research topics of neuromorphic computing to generative tasks and applications.
6 Acknowledgment
This work was supported in part by CoCoSys, a JUMP2.0 center sponsored by DARPA and SRC, the National Science Foundation (CAREER Award, Grant #2312366, Grant #2318152), the DARPA Young Faculty Award and the DoE MMICC center SEA-CROGS (Award #DE-SC0023198).